2-(50) Find Is, Ic , vce and the region of operation, 20012 IM? HI 3v AW 200 6- 200
Answer #1

Tov 2001 To f?-200 2- LE IMV I AL 3= 200 (183+.Ie) - IMEB + YE VE The grob VG 521911 V 3= 200 (IB PB Tr] +1M57+08 F.3 = 2001